Find values for m and b in the following system so that the solution to the system is (-3,4).
5x + 7y = b
mx + y = 22

Answer by praseenakos@yahoo.com(507)

QUESTION:

Find values for m and b in the following system so that the solution to the system is (-3,4).
5x + 7y = b
mx + y = 22


ANSWER:



Solution of a system of equation means the set of values which satisfies the
system of equations.




Here the system of equations are in two variables, that is in x and y.




Now substitute these values x=-3 and y = 4 in the given system of equations.



Then the equations become,



5*-3 + 7*4 = b

==> -15 + 28 = b

==> 13 = b





m*-3 + 4 = 22


==> -3m + 4 = 22


==> -3m = 22-4


==> -3m = 18


==> m = 18/-3




==> m = -5


So the values are,

m = -5
b = 13



Hope you understood.


Regards.


praseenakos@yahoo.co.in